OpenStudy (anonymous):

PLEASEEE HELP! Graph each circle so that the center is at the origin. Then write the equation. radius 6

OpenStudy (anonymous):

center at the origin make it look like \(x^2+y^2=r^2\) in your case \(r=6\)

OpenStudy (anonymous):

how do I graph the circle? @satellite73